OpenStudy (anonymous):

A property owner wants to fence a garden pot adjacent to a road. The fencing next to the road must be sturdier and costs $5 per foot, but the other fencing costs just $3 per foot. The garden is to have an area of 1200 square feet. Find a function that models the cost of fencing the garden, and use this function to find the garden dimensions that minimizes the cost of fencing.

OpenStudy (anonymous):

put the part next to the road as x, and the adjacent side as y as in the picture. then you know that \[xy=1200\] meaning \[y=\frac{1200}{x}\]

OpenStudy (anonymous):

meant \[xy=1200\], \[y=\frac{1200}{x}\]

OpenStudy (anonymous):

cost is \[C(x)=5x+3x+3\times \frac{1200}{x}+3\times \frac{1200}{x}=8x+\frac{7200}{x}\] that is what you want to minimize

OpenStudy (anonymous):

take the derivative, set it equal zero to find the critical points, solve for x and you are done